Assistance with a formula to find the average of 35%, 20%, 15%, 25%, and 5%.
I am a transitional student, and need lots of help in this area. I thank you in advance for your assistance.

Ms. Daphne Roach

Answer by Fombitz(13823)

To find an average, sum all of the points and divide by the number of points.
Sum= 35+20+15+25+5=100
Number of points = 5
Average = Sum/Number =100/5=20
The average is 20%.